OpenStudy (anonymous):

solve by clearing fractions. -2 + 3/4x=5x - 1/3 one more thing, if a problem has fractions dont you always solve it by clearing the fractions first?

myininaya (myininaya):

\[-2+\frac{3}{4}x=5x-\frac{1}{3}\] multiply both sides by 12 \[-24+9x=60x-4\] \[-24+4=60x-9x\] \[-20=51x\] \[\frac{-20}{51}=x\] --------------------------------- OR: \[-2+\frac{1}{3}=5x-\frac{3}{4}x\] find commom denominator \[\frac{-2*3}{3}+\frac{1}{3}=\frac{5*4}{4}x-\frac{3}{4}x\] \[\frac{-6+1}{3}=\frac{20-3}{4}x\] \[\frac{-5}{3}=\frac{17}{4}x\] \[\frac{-5*4}{3*17}=x\] \[x=\frac{-20}{51}\]